Grade 12th passAlgebra

Find the square of the following numbers.
32 …......

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

(32)2
= (30 +2)2
= (30)2 + (2)2 + 2×30×2 [Since, (a+b)2 = a2+b2 +2ab]
=900+4+120
= 1024

Thanks